# DRUPI: Dataset Reduction Using Privileged Information

### v2

@@ 3.3 Learning Using Privileged Information
We have previously discussed how to synthesize appropriate privileged information $f_{i}^{\star}$ for a given reduced dataset $\mathcal{D}_{\mathcal{S}}$ and extend it into $\mathcal{D}_{\mathcal{S}}^{\star}$. We now focus on leveraging the new reduced dataset $\mathcal{D}_{\mathcal{S}}^{\star}$ to enhance a model’s performance on unseen test data. Following the learning using privileged information (LUPI) framework (Pechyony & Vapnik, 2010; Lopez-Paz et al., 2015), we incorporate the additional privileged information $f_{i}^{\star}$ during training to build a classifier that outperforms those trained solely on the regular reduced dataset $\mathcal{D}_{\mathcal{S}}$.
@@
3.4 Theoretical Analysis
We propose a theoretical analysis to elucidate the mechanisms by which the DRUPI framework enhances the quality of reduced datasets. Let $g_{\mathcal{T}}\in\mathcal{G}_{\mathcal{T}}$ denote the oracle function for the original dataset $\mathcal{D}_{\mathcal{T}}$, and $|\cdot|_{C}$ represent a function class capacity measure, i.e., a measure of model performance. Consider two models: one trained on the pure reduced dataset $\mathcal{D}_{\mathcal{S}}$, represented by $g_{\mathcal{S}}\in\mathcal{G}_{\mathcal{S}}$, and another trained on the reduced dataset with privileged information, $\mathcal{D}_{\mathcal{S}}^{\star}$, denoted by $g_{\mathcal{S}^{\star}}\in\mathcal{G}_{\mathcal{S}^{\star}}$.
We begin by briefly reviewing the well-known VC theory (Vapnik, 1998), a fundamental analytical tool in statistical learning theory that forms the basis of our theoretical framework. The VC-dimension defines the performance of a model to limited data points. Specifically, for a model $g$ belonging to a function class $\mathcal{G}$, with a finite VC-dimension $|\mathcal{G}|_{\text{VC}}$, the expected error $R(g)$ with probability $1-\delta$ is bounded as follows:
where the $O(\cdot)$ term is the estimation error, and $R_{m}(g)$ is the training error over $m$ data points, and $\alpha$ lies between $\frac{1}{2}$ and $1$. The parameter $\alpha$ represents the difficulty of the task. For more difficult, non-separable tasks, $\alpha\approx\frac{1}{2}$, yielding a slower learning rate of $O(m^{-1/2})$. For easier, separable tasks, where the model makes no training errors, $\alpha\approx 1$, yielding a faster learning rate of $O(m^{-1})$. Given a student learning from a fixed dataset of size $m$, a good teacher model can ease the learning process by accelerating the learning rate from $O(m^{-1/2})$ to $O(m^{-1})$.
Next, we provide theoretical analysis for DRUPI, showing how incorporating privileged information accelerates the learning process and improves the quality of the reduced dataset. Building on the top of (Lopez-Paz et al., 2015), we extend the results in dataset reduction scenarios. First, assume that the model trained on the pure reduced dataset $g_{\mathcal{S}}$ learns the true function $g_{\mathcal{T}}$ at a slower rate $\alpha_{\mathcal{S}}$:
where the final inequality arises because $\alpha\leq 1$. Therefore, for $\frac{1}{2}<\alpha\leq 1$ in dataset reduction settings, the inequality becomes:
This inequality highlights the advantages of models trained with privileged information: training using privileged information exhibit lower generalization and approximation errors compared to those trained without privileged information. More importantly, it emphasizes that the privileged information is most beneficial in low-data regimes, which is the typical DR scenario. These benefits align with the principles of LUPI as outlined in  (Vapnik et al., 2015; Lopez-Paz et al., 2015).
@@ 4 Experiments
Table 1: Application of DRUPI to representative pruning methods on CIFAR-10/100. We initialized the dataset with the baseline methods, and utilized DC for synthesizing feature labels.
@@
Random 13.5$\pm$0.4 20.0$\pm$0.5 27.1$\pm$0.6 35.8$\pm$0.6 43.0$\pm$0.5 4.3$\pm$0.2 9.5$\pm$0.2 14.5$\pm$0.2 19.5$\pm$0.4 29.5$\pm$0.3
L-Conf 10.7$\pm$0.4 10.5$\pm$0.4 10.8$\pm$0.4 17.9$\pm$0.4 23.1$\pm$0.6 2.1$\pm$0.1 3.6$\pm$0.1 6.6$\pm$0.2 9.0$\pm$0.2 16.4$\pm$0.3
Entropy 12.2$\pm$0.6 14.1$\pm$0.5 14.8$\pm$0.5 19.6$\pm$0.4 23.8$\pm$0.7 1.7$\pm$0.1 3.7$\pm$0.2 6.7$\pm$0.2 9.0$\pm$0.3 17.1$\pm$0.3
Margin 8.9$\pm$0.4 15.8$\pm$0.7 20.3$\pm$0.4 24.8$\pm$0.5 31.3$\pm$0.5 3.0$\pm$0.2 6.2$\pm$0.2 9.0$\pm$0.2 12.7$\pm$0.3 20.7$\pm$0.3
Glister 11.5$\pm$0.3 16.9$\pm$0.5 23.0$\pm$0.4 28.4$\pm$0.5 30.0$\pm$0.5 2.9$\pm$0.5 7.1$\pm$0.4 10.4$\pm$0.6 13.3$\pm$0.7 27.2$\pm$0.7
Graig 18.1$\pm$0.3 19.5$\pm$0.4 19.0$\pm$0.5 27.8$\pm$0.3 30.2$\pm$0.4 4.3$\pm$0.5 9.0$\pm$0.4 13.6$\pm$0.7 14.6$\pm$0.5 20.1$\pm$0.6
Herding 15.3$\pm$0.5 23.5$\pm$0.3 25.1$\pm$0.5 26.7$\pm$0.4 34.9$\pm$0.6 4.0$\pm$0.1 6.2$\pm$0.2 8.1$\pm$0.4 13.1$\pm$0.6 18.4$\pm$0.6
+DRUPI 27.9$\pm$0.6 37.3$\pm$0.6 45.8$\pm$0.7 51.0$\pm$0.4 54.0$\pm$0.6 14.0$\pm$0.3 20.4$\pm$0.5 25.5$\pm$0.4 28.9$\pm$0.3 31.4$\pm$0.6
12.6 13.8 20.7 24.3 19.1 10.0 14.2 17.4 15.8 13.0
k-Center 16.4$\pm$0.6 22.4$\pm$0.6 23.1$\pm$0.5 30.4$\pm$0.4 36.7$\pm$0.5 4.8$\pm$0.2 6.7$\pm$0.4 10.0$\pm$0.5 15.9$\pm$1.1 21.8$\pm$1.0
+DRUPI 29.7$\pm$0.6 40.0$\pm$0.6 46.2$\pm$0.6 50.8$\pm$0.6 54.3$\pm$0.6 13.5$\pm$0.3 20.0$\pm$0.5 25.9$\pm$0.3 29.1$\pm$0.5 32.0$\pm$0.5
13.3 17.6 23.1 20.4 17.6 8.7 13.3 15.9 13.2 10.2
Forgetting 15.3$\pm$0.6 19.1$\pm$0.7 23.9$\pm$0.7 26.9$\pm$0.7 39.5$\pm$0.5 4.1$\pm$0.1 7.8$\pm$0.3 10.4$\pm$0.5 14.1$\pm$0.6 22.3$\pm$0.4
+DRUPI 30.0$\pm$0.6 39.7$\pm$0.7 46.6$\pm$0.6 51.3$\pm$0.5 54.5$\pm$0.5 14.0$\pm$0.4 20.1$\pm$0.6 25.8$\pm$0.3 29.3$\pm$0.4 32.2$\pm$0.5
14.7 20.6 22.7 24.4 15.0 9.9 12.3 15.4 15.2 9.9
Full Dataset 84.8$\pm$0.1 56.2$\pm$0.3
Table 2: Results of DRUPI on distillation methods across CIFAR-10/100, and Tiny ImageNet. We initialized reduced datasets with corresponding baseline methods, and synthesized feature labels for these datasets with the same baseline methods.
@@
IPC 1 10 50 1 10 50 1 10
Random 15.4$\pm$0.3 31.0$\pm$0.5 50.6$\pm$0.3 4.2$\pm$0.3 14.6$\pm$0.5 33.4$\pm$0.4 1.4$\pm$0.1 5.0$\pm$0.2
KIP 49.9$\pm$0.2 62.7$\pm$0.3 68.6$\pm$0.3 15.7$\pm$0.2 28.3$\pm$0.1 - - -
DM 26.0$\pm$0.8 48.9$\pm$0.6 63.0$\pm$0.4 11.4$\pm$0.3 29.7$\pm$0.3 43.6$\pm$0.4 3.9$\pm$0.2 12.9$\pm$0.4
DSA 28.8$\pm$0.7 52.1$\pm$0.5 60.6$\pm$0.5 13.9$\pm$0.3 32.3$\pm$0.3 42.8$\pm$0.4 - -
DCC 32.9$\pm$0.8 49.4$\pm$0.5 61.6$\pm$0.4 13.3$\pm$0.3 30.6$\pm$0.4 40.0$\pm$0.3 - -
DSAC 34.0$\pm$0.7 54.5$\pm$0.5 64.2$\pm$0.4 14.6$\pm$0.3 33.5$\pm$0.3 39.3$\pm$0.4 - -
CAFE 30.3$\pm$1.1 46.3$\pm$0.6 55.5$\pm$0.6 12.9$\pm$0.3 27.8$\pm$0.3 37.9$\pm$0.3 - -
IDM 45.6$\pm$0.7 58.6$\pm$0.1 67.5$\pm$0.1 20.1$\pm$0.3 45.1$\pm$0.1 50.0$\pm$0.2 - -
DC 28.3$\pm$0.5 44.9$\pm$0.5 53.9$\pm$0.5 12.8$\pm$0.3 25.2$\pm$0.3 29.8$\pm$0.3 - -
+DRUPI 31.5$\pm$0.9 47.4$\pm$0.9 55.0$\pm$0.5 14.9$\pm$0.4 29.2$\pm$0.5 30.9$\pm$0.5 - -
3.2 2.5 1.1 2.1 4.0 1.1
MTT 46.2$\pm$0.8 65.4$\pm$0.7 71.6$\pm$0.2 24.3$\pm$0.3 39.7$\pm$0.4 47.7$\pm$0.2 8.8$\pm$0.3 23.2$\pm$0.2
+DRUPI 47.4$\pm$0.5 65.8$\pm$0.6 71.7$\pm$0.2 25.6$\pm$0.4 40.8$\pm$0.3 48.8$\pm$0.3 11.2$\pm$0.1 24.9$\pm$0.2
1.2 0.4 0.1 1.3 1.1 1.1 2.4 1.7
Full Dataset 84.8$\pm$0.1 56.2$\pm$0.3 37.6$\pm$0.4
@@ 4.1 Experimental Setup
In this section, we investigate the effectiveness of our proposed method, DRUPI, through a series of experiments on diverse datasets and tasks. We begin by evaluating the efficacy of DRUPI when applied to coreset selection and dataset distillation tasks. Specifically, we followed prior works to conduct experiments on CIFAR-10/100 (Krizhevsky et al., 2009) for coreset selection methods, where ResNet-18 (He et al., 2016) is utilized for extracting importance score. For the dataset distillation methods, we conducted experiments on CIFAR-10/100, Tiny ImageNet (Le & Yang, 2015), and subsets of ImageNet (Russakovsky et al., 2015).
For coreset selection, we benchmarked against several representative baselines, including Random, L-conf, Entropy, Margin (Coleman et al., 2019a), Glister (Killamsetty et al., 2021b), Graig (Mirzasoleiman et al., 2020), Herding (Welling, 2009), k-Center (Har-Peled & Mazumdar, 2004), and Forgetting (Toneva et al., 2018). More detailed settings are provided in Appendix A.2.
For dataset distillation, we evaluated a range of advanced methods, including KIP (Nguyen et al., 2020), DM (Zhao & Bilen, 2022), DSA (Zhao & Bilen, 2021), DCC, DSAC (Lee et al., 2022), CAFE (Wang et al., 2022), IDM (Zhao et al., 2023), DC (Zhao et al., 2020), and MTT (Cazenavette et al., 2022). In line with prior studies, we used networks with instance normalization as the default setting. Unless otherwise specified, distillation was performed with a depth-3 ConvNet for CIFAR-10/100, a depth-4 ConvNet for Tiny ImageNet, and a depth-5 ConvNet for ImageNet subsets. See Appendix A.3 for more details.
It is worth noting that for both pruning and distillation methods, we initialized all data-label pairs using the baseline method and employed a weakly-trained model (trained for one epoch) to extract feature labels, which were then used to synthesize privileged information. By default, we utilized DC to synthesize one feature label per data-label pair, aligning it with features extracted from the final layer of a ConvNet during bi-level optimization. We employed $\lambda_{reg}=0.5$ and $\lambda_{task}=0.1$, and learning rate = 0.1 for feature optimization.
@@ 4.2 Main Results
Coreset selection. In our experiments, DRUPI utilizes the reduced dataset initialized with the Herding, k-Center, and Forgetting methods to assess its performance across diverse fraction on CIFAR-10/100. As shown in Table 1, by incorporating privileged information, these methods consistently outperformed the baseline across a range of fraction settings on CIFAR-10/100. Particularly, on the CIFAR-10 (fraction = 0.4%) , DRUPI achieved a performance increase of 24.4% on the Forgetting method and 24.3% on the Herding method. We find that for datasets without optimized instances (e.g., selected coresets), the performance gain is much higher than those with optimized samples.
Dataset distillation. We employed DRUPI in several classical dataset distillation methods, where privileged information is obtained with DC and MTT. Table 2 summarizes the classification performances of ConvNets trained with different distillation methods. Specifically, applying DRUPI to DC on CIFAR-100 (10 IPC) resulted in a 4% improvement. For MTT, DRUPI delivered a 2.4% gain on Tiny ImageNet (1 IPC). Additionally, we evaluated its effectiveness on ImageNet subsets, as shown in Table 3, where DRUPI applied to MTT led to a 3.4% improvement on ImageMeow with 10 IPC, demonstrating strong performance even on larger datasets. Notably, learning both feature labels with DRUPI outperforms simply extracting features alone. Further results for DATM are provided in Appendix B.1.
Table 3: Results on ImageNet subsets when integrating DRUPI into dataset distillation methods. Reduced datasets are initialized with MTT, and feature labels are synthesized with MTT.
@@
IPC 1 10 1 10 1 10 1 10 1 10
MTT 47.7$\pm$0.9 63.0$\pm$1.3 28.6$\pm$0.8 35.8$\pm$1.8 26.6$\pm$0.8 40.3$\pm$1.3 30.7$\pm$1.6 40.4$\pm$2.2 45.2$\pm$0.8 60.0$\pm$1.5
+DRUPI 50.5$\pm$0.1 65.7$\pm$0.5 31.3$\pm$0.2 37.5$\pm$1.0 29.1$\pm$1.4 43.0$\pm$0.9 34.0$\pm$1.6 43.8$\pm$0.9 46.6$\pm$0.6 62.2$\pm$0.9
2.8 2.7 2.7 2.7 2.5 2.7 3.3 3.4 1.4 2.2
@@
Table 5: Cross-architecture evaluations of distilled datasets on unseen networks. Reduced datasets are initialized with DC on CIFAR-10 (IPC=10). Feature labels are learned with DC.
DC +DRUPI DC +DRUPI DC +DRUPI
@@
LeNet 34.0$\pm$0.7 34.9$\pm$0.7 (0.9) 32.6$\pm$0.5 34.3$\pm$0.5 (1.7) 25.6$\pm$2.1 32.4$\pm$2.9 (6.8)
MLP 22.3$\pm$1.0 34.6$\pm$0.7 (12.3) 36.6$\pm$0.5 37.1$\pm$0.4 (0.5) 29.9$\pm$2.5 30.0$\pm$0.9 (0.1)
ResNet 31.6$\pm$0.9 32.6$\pm$0.7 (1.0) 28.3$\pm$0.9 30.0$\pm$0.6 (1.7) 24.4$\pm$3.3 27.2$\pm$3.6 (2.8)
VGG 30.1$\pm$0.7 34.5$\pm$0.9 (4.4) 27.0$\pm$0.7 31.8$\pm$1.1 (4.8) 22.5$\pm$2.5 27.1$\pm$3.2 (4.6)
ConvNet 35.4$\pm$0.6 35.8$\pm$0.8 (0.4) 28.0$\pm$0.8 32.3$\pm$1.0 (4.3) 20.1$\pm$4.5 27.3$\pm$3.2 (7.2)
AlexNet 33.8$\pm$0.7 34.3$\pm$0.8 (0.5) 31.0$\pm$0.6 32.3$\pm$0.5 (1.3) 34.0$\pm$4.6 38.2$\pm$1.7 (4.2)
4.3 Cross-architecture generalization
Cross-architecture evaluation is a critical step toward ensuring robust generalization across previously unseen architectures. We measured the quality of reduced dataset with privileged information on both pruning and distillation settings. To address the misalignment between the shapes of the learned feature/attention labels and the intermediate features of different network architectures, we introduced an additional fully connected layer that is trained alongside the evaluation model.
For distillation methods, we applied DRUPI to synthesize feature labels for reduced datasets initialized by DC and MTT. As detailed in Table 5, we conducted experiments on distilled dataset initialized by DC on CIFAR-10 (10 IPC), and trained models on 6 distinct network architectures, and evaluated the model performance across them. Notably, DRUPI achieved an 18.3% performance improvement over the baseline when training on VGG and evaluating on ResNet. Additional results on attention labels are provided in Appendix B.2. Table 9 presents the results of applying DRUPI to MTT on CIFAR-10 (1 IPC), showing significant gains, such as an 11.1% improvement when training on ConvNet and evaluating on AlexNet.

### v3

@@ 3.3 Learning Using Privileged Information
As discussed above, we have discussed how to synthesize appropriate privileged information $f_{i}^{\star}$ for a given reduced dataset $\mathcal{D}_{\mathcal{S}}$ and extend it into $\mathcal{D}_{\mathcal{S}}^{\star}$. We now focus on leveraging the new reduced dataset $\mathcal{D}_{\mathcal{S}}^{\star}$ to enhance a model’s performance on unseen test data. Following the learning using privileged information (LUPI) framework Pechyony and Vapnik (2010); Lopez-Paz et al. (2015), we incorporate the additional privileged information $f_{i}^{\star}$ during training to build a classifier that outperforms those trained solely on the regular reduced dataset $\mathcal{D}_{\mathcal{S}}$.
@@ 4 Experiments
Table 1: Application of DCPI to representative pruning methods on CIFAR-10/100. We initialized the dataset using the baseline methods and utilized DC for synthesizing feature labels, with the same memory cost for fair compairason.
@@
[…]
Table 2: Results of DCPI on distillation methods across CIFAR-10/100, and Tiny ImageNet. We initialized reduced datasets with corresponding baseline methods, and synthesized feature labels for these datasets with the same baseline methods.
@@
Fraction (%) 0.02 0.2 1 0.2 2 10 0.2 2
[…]
@@ 4.1 Experimental Setup
In this section, we investigate the effectiveness of our proposed method, DCPI, through a series of experiments on diverse datasets and tasks. We begin by evaluating the efficacy of DCPI when applied to coreset selection and dataset distillation tasks. Specifically, we followed prior works to conduct experiments on CIFAR-10/100 Krizhevsky et al. (2009) for coreset selection methods, where ResNet-18 He et al. (2016) is utilized for extracting importance score. For the dataset distillation methods, we conducted experiments on CIFAR-10/100, Tiny ImageNet Le and Yang (2015), and subsets of ImageNet Russakovsky et al. (2015).
For coreset selection, we benchmarked our method against several representative baselines, including Random, L-conf, Entropy, Margin Coleman et al. (2019), Glister Killamsetty et al. (2021b), Graig Mirzasoleiman et al. (2020), Herding Welling (2009), k-Center Har-Peled and Mazumdar (2004), and Forgetting Toneva et al. (2018). More detailed settings are provided in Appendix A.2.
For dataset distillation, we evaluated advanced methods, including KIP Nguyen et al. (2020), DM Zhao and Bilen (2022), DSA Zhao and Bilen (2021), DCC Lee et al. (2022b), DSAC Lee et al. (2022a), CAFE Wang et al. (2022), IDM Zhao et al. (2023), DC Zhao et al. (2020), and MTT Cazenavette et al. (2022). In line with prior studies, we used networks with instance normalization as the default setting. Unless otherwise specified, distillation was performed with a depth-3 ConvNet for CIFAR-10/100, a depth-4 ConvNet for Tiny ImageNet, and a depth-5 ConvNet for ImageNet subsets. Table 8 and Table 9 in Appendix A.3 provided detailed information regarding hyperparameter settings.
It is worth noting that for both pruning and distillation methods, we initialized all data-label pairs using the baseline method and employed a weakly-trained model (trained for one epoch) to extract feature labels, which were then used to synthesize privileged information. By default, we utilized DC to synthesize one feature label per data-label pair, aligning it with features extracted from the final layer of a ConvNet during bi-level optimization.
@@ 4.2 Main Results
Coreset selection. In our experiments, DCPI utilizes the reduced dataset initialized with the Herding, k-Center, and Forgetting methods to assess its performance across diverse fraction on CIFAR-10/100. To ensure a fair comparison, our experiments in this work are conducted under the same memory constraints as the baseline methods. As shown in Table 1, by incorporating privileged information, these methods consistently outperformed the baseline across a range of fraction settings on CIFAR-10/100. Particularly, on the CIFAR-10 (fraction = 0.4%) , DCPI achieved a performance increase of 24.4% on the Forgetting method and 24.3% on the Herding method. We find that for datasets without optimized instances (e.g., selected coresets), the performance gain is much higher than that with optimized samples.
Dataset distillation. We employed DCPI in several methods, where privileged information is obtained with DC and MTT and RDED. Table 2 summarizes the classification performances of ConvNets trained with different distillation methods. Specifically, applying DCPI to DC on CIFAR-100 (fraction = 0.2%) resulted in a 2.1% improvement. For MTT, DCPI delivered a 2.4% gain on Tiny ImageNet (fraction = 0.2%). Additionally, we evaluated its effectiveness on ImageNet subsets, as shown in Table 3, where DCPI applied to MTT led to a 3.4% improvement on ImageMeow with fraction = 1%, demonstrating strong performance even on larger datasets. Notably, learning both feature labels with DCPI outperforms simply extracting features alone. Further results for DATM are provided in Appendix B.1. The application of DCPI to RDED is detailed in Tables 2 and 4. Specifically, DCPI achieved a 12.9% improvement on CIFAR-100 (fraction = 0.2%). Furthermore, its effectiveness extends to large-scale datasets like ImageNet-1k. For instance, when applied with fraction = 0.08%, DCPI improved the classification performance of ResNet-18 by 4.6%. Additionally, for ResNet-101 with fraction = 3.90%, DCPI yielded a 2.9% performance gain, highlighting its robustness across different model architectures and dataset scales.
Table 3: Results on ImageNet subsets when integrating DCPI into dataset distillation methods. Reduced datasets are initialized with MTT, and feature labels are synthesized with MTT.
@@
fraction(%) 0.1 1 0.1 1 0.1 1 0.1 1 0.1 1
[…]
Table 4: Performance on ImageNet-1K when integrating DCPI into dataset distillation methods. Reduced datasets are initialized with RDED, and feature labels are synthesized with RDED. Data was distilled using both ResNet-18 and ConvNet, and the same models were employed for evaluation. In addition, we validated the dataset distilled by ResNet-18 using ResNet-101.
Fraction (%) ResNet-18 ResNet-101 ConvNet
SRe2L RDED +DCPI SRe2L RDED +DCPI MTT IDM TESLA DATM RDED +DCPI
[…]
@@
4.3 Cross-architecture generalization
Cross-architecture evaluation is a critical step toward ensuring robust generalization across previously unseen architectures. We measured the quality of reduced dataset with privileged information on both pruning and distillation settings. Incorporating diverse privileged information lets us tailor datasets to specific needs, while an additional fully connected layer resolves shape misalignment and maximizes reduced data potential.
For distillation methods, we applied DCPI to synthesize feature labels for reduced datasets initialized by DC and MTT. As detailed in Table 6 presents the results of applying DCPI to MTT on CIFAR-10 (fraction = 0.02%), showing significant gains, such as an 11.1% improvement when training on ConvNet and evaluating on AlexNet. Table 7, we conducted experiments on distilled dataset initialized by DC on CIFAR-10 (fraction = 0.2%), and trained models on 6 distinct network architectures, and evaluated the model performance across them. Notably, DCPI achieved an 18.3% performance improvement over the baseline when training on VGG and evaluating on ResNet.
In addition to the results presented in Table 5 and 7, which summarize the performance of utilizing the synthesized datasets from pruning and dataset distillation to initialize the feature labels, we provide additional results in Table 6 to evaluate the cross-architecture generalization by using both feature labels and attention labels on different reduced datasets.
@@ 4.3 Cross-architecture generalization
[…]
Table 6 illustrates the cross-architecture evaluations conducted for both dataset condensation (DC) and MTT across various network architectures. In these experiments, we propose DCPI-F (feature labels learned by DC) and its variant DCPI-A, which is further enhanced by pooling through spatial attention. Similarly, feature labels synthesized with MTT are also average pooled into channel attention labels. We utilized several networks for cross evaluation, including LeNet, MLP, ResNet, VGG, ConvNet, and AlexNet.
The table demonstrates that initializing the reduced dataset with feature labels assigned through intermediate features of pre-trained networks leads to significant performance improvements across all architectures. Moreover, further gains are observed when channel attention pooling is applied. By leveraging Eq. (5) to learn and update the feature labels, the reduced dataset consistently yields competitive or improved performance across different networks. This highlights the benefits of feature and attention labels in enhancing model generalization, with particularly notable improvements observed in ConvNet and AlexNet settings. Channel attention pooling further contributes to these gains, reinforcing the effectiveness of this approach across architectures.
Table 7: Cross-architecture evaluations of distilled datasets on unseen networks. Reduced datasets are initialized with DC on CIFAR-10 (fraction = 0.2%). Feature labels are learned with DC.
[…]
DC +DCPI DC +DCPI DC +DCPI
[…]